### 13 Mar 2026 v3

# Euclid: A blue galaxy population and a brightest cluster galaxy in the making in a MaDCoWS2 galaxy cluster candidate††thanks: This paper is published on behalf of the Euclid Consortium

Euclid: A blue galaxy population and a brightest cluster galaxy in the making in a MaDCoWS2 galaxy cluster candidate††thanks: This paper is published on behalf of the Euclid Consortium
We present an example cluster follow-up study with Euclid. Our target, a candidate cluster nicknamed the ‘Puddle’, was initially discovered by the Massive and Distant Clusters of WISE Survey 2 as a candidate cluster. It was also detected independently as a candidate with the two cluster-finding algorithms in Euclid Quick Release 1 (Q1). A Keck MOSFIRE spectrum shows the brightest nucleus is at and is dominated by an active galactic nucleus. Our analysis focused on the galaxy population and the brightest cluster galaxy (BCG), and is based on Euclid and ancillary photometry. Compared to similar fields, we measured an overdensity of galaxies with in a radius around the BCG. About % of the completeness-corrected galaxy population is red, which is consistent with some clusters at but lower than others. Euclid imaging revealed that six or seven galaxies appear to be assembling to form the future BCG. Spectral energy distribution fitting suggests that the merging BCG has a stellar mass of and that it experienced a short burst of star formation Myr ago. Its morphology and star-formation history suggest that the proto-BCG is a more evolved version of the merging core of SPT234956. These systems indicate that multiobject mergers might be a common BCG formation process. Assuming a similar density of mergers in the Euclid Wide Survey, we expect that Euclid will discover approximately 400 assembling BCGs by the end of its mission.
Key Words.:
Galaxies: active, Galaxies: clusters: individual: MOO2 J0337428386/EUCLQ1CL J033730.18283827.6, Galaxies: formation, Galaxies: interactions, Galaxies: star formation1 Introduction
Galaxy clusters sit at the nodes of the cosmic web and are linked by filaments (Bond et al. 1996; Springel et al. 2005; Vogelsberger et al. 2014). As the maxima of the density field in the Universe, galaxy clusters are natural laboratories to test a variety of astrophysical phenomena in different fields: cosmology (e.g. Vikhlinin et al. 2009; Allen et al. 2011; Pierre et al. 2016), plasma and high-energy physics (e.g. McNamara and Nulsen 2007, 2012; Zhuravleva et al. 2014), and galaxy evolution (e.g. von der Linden et al. 2010; Peng et al. 2010b, 2012).
The cessation of star formation in galaxies, usually called ‘quenching’, depends on two main factors: galaxy mass (Kauffmann et al. 2003) and environment density (Peng et al. 2010b, 2012). Dense environments such as local galaxy clusters tend to host a larger fraction of quiescent galaxies than field galaxy samples (e.g. Dressler 1980; Poggianti et al. 1999; Balogh et al. 2004, 2009, 2016; Kawinwanichakij et al. 2017; Pintos-Castro et al. 2019; Ragusa et al. 2025). At the fraction of quenched galaxies in clusters decreases with increasing redshift (e.g. Balogh et al. 2004; Raichoor and Andreon 2012; Pintos-Castro et al. 2019) but always remains above the field level. The situation is less clear at . Some authors find that high-redshift galaxy clusters already exhibit elevated levels of quenching (e.g. Newman et al. 2014; Cooke et al. 2019; Lemaux et al. 2019; Strazzullo et al. 2019; van der Burg et al. 2020; Toni et al. 2026), while others observe no difference with field levels before (e.g. Tran et al. 2010; Brodwin et al. 2013; Alberts et al. 2016; Nantais et al. 2020; Trudeau et al. 2024).
Brightest cluster galaxies (BCGs) represent another example of the interplay between galaxy evolution and cluster environment. BCGs are the most massive galaxies in the Universe, and they are found close to the centre of the potential wells of galaxy clusters (Zitrin et al. 2012; Hashimoto et al. 2014; Cui et al. 2016; Lopes et al. 2018). Their properties (e.g. shape, mass, star-formation history) tend to correlate with those of their host clusters (e.g. Ebeling et al. 2021). A particular example is the relation between the active galactic nucleus (AGN) activity in BCGs and the regulation of the temperature of the intracluster medium (ICM; Hu et al. 1985; Burns 1990; Cavagnolo et al. 2008; Rafferty et al. 2008; Hlavacek-Larrondo et al. 2012). An imbalance between the cooling of the ICM and the energy injected by the AGN (e.g. McNamara and O’Connell 1989; Fabian 1994; McNamara et al. 2005; Rafferty et al. 2006) can trigger gas condensation onto the BCG, a phenomenon called a ‘cooling flow’.
The BCG evolution is intimately linked to the evolution of its host cluster. At , BCG evolution is dominated by gas-poor mergers (e.g. Aragón-Salamanca et al. 1998; Dubinski 1998; Lidman et al. 2012, 2013; Burke and Collins 2013; Golden-Marx and Miller 2018; Golden-Marx et al. 2025; Montenegro-Taborda et al. 2023). While most local BCGs are quiescent, between 20% and 35% of them exhibit small amounts of star formation (usually below the field level, see Orellana-González et al. 2022) powered by residual cooling flows (e.g. Crawford et al. 1999; Rawle et al. 2012; Oliva-Altamirano et al. 2014; McDonald et al. 2016).
However, the formation and early evolution of BCGs remain poorly understood. A popular formation model (De Lucia and Blaizot 2007) predicted that 80% of the stellar mass of BCGs would form before in separate progenitors that progressively assemble into a BCG through gas-poor mergers. Observations of BCGs have shown a different picture: High-redshift BCGs are diverse, but some of them display substantial levels of in situ star formation (Webb et al. 2015a; McDonald et al. 2016; Bonaventura et al. 2017). The cause of the star formation is not currently understood. Some authors (e.g. Rennehan et al. 2020; Remus et al. 2023) suggest that it might be powered by the initial formation of the BCG in a gas-rich merger involving multiple galaxies. Observations of protocluster cores with elevated star-formation rates (SFRs) and interacting galaxies support this scenario (Miley et al. 2006; Kuiper et al. 2011; Miller et al. 2018; Coogan et al. 2023). One cluster core at , SpARCS104922.6+564032.5 (Webb et al. 2015b; Hlavacek-Larrondo et al. 2020) indicates that massive cooling flows could also power star formation.
Our current understanding of quenching in high-redshift clusters and BCG early evolution are limited by small number statistics since most cluster samples contain only a few objects at (e.g. Wilson et al. 2006, 2009; Wylezalek et al. 2013; Bleem et al. 2015; Pierre et al. 2016; Bulbul et al. 2022, 2024). There are several recent protocluster searches that explore these redshifts (e.g. Martinache et al. 2018; Ouchi et al. 2018; Toshikawa et al. 2018; Golden-Marx et al. 2019; Ando et al. 2020; Gully et al. 2024), but their sample purities tend to be low (e.g. Hung et al. 2025).
With the advent of a new generation of large-scale survey facilities (Euclid, Nancy Grace Roman, Vera C. Rubin, see Spergel et al. 2015; Ivezić et al. 2019; Robertson et al. 2019; Euclid Collaboration: Scaramella et al. 2022; Grishin et al. 2025), this situation is changing, and large samples of (proto)clusters are becoming available. Euclid is the only one of these facilities that is currently in service. Located at the second Sun-Earth Lagrange point, Euclid has a 1.2-metre primary mirror (Euclid Collaboration: Mellier et al. 2025). By the end of its six-year mission, Euclid is expected to observe deg2 of the sky (Euclid Collaboration: Mellier et al. 2025) and to find about galaxy clusters at , according to the latest estimates (Euclid Collaboration: Mellier et al. 2025, see also the estimations of Euclid Collaboration: Adam et al. 2019 and Sartoris et al. 2016).
The Euclid search for galaxy clusters uses two algorithms: the Adaptive Matched Identifier of Clustered Objects (AMICO) and PZWav (Euclid Collaboration: Adam et al. 2019; Euclid Collaboration: Bhargava et al. 2025). AMICO is a matched filter algorithm and it detects clusters with a redshift-dependent filter mimicking the sum of the cluster contribution and the noise (Bellagamba et al. 2018; Maturi et al. 2019). PZWav instead applies a spatial filter constructed from the difference of two Gaussians. The width of the Gaussians was chosen to select overdensities with sizes consistent with a cluster extent on the sky (Thongkham et al. 2024b).
The PZWav algorithm is also used by another large-scale cluster survey, the Massive and Distant Clusters of WISE Survey 2 (MaDCoWS2; Thongkham et al. 2024b, a). MaDCoWS2 derives photometric redshifts (Brodwin et al. 2006) using data from the Dark Energy Camera (DECam) and the Wide-Field Infrared Survey Explorer (WISE), and it uses these photometric redshifts as input to PZWav. While the MaDCoWS2 survey identified 6959 candidate clusters at , the resolution of WISE – the point spread functions (PSFs) full widths at half maxima (FWHM) are and in the two bluest channels (see Wright et al. 2010) – is insufficient for distinguishing individual galaxies in the packed cores of high-redshift clusters. Near-infrared data sets with better resolution, such as Euclid images, are needed to enable galaxy evolution studies of MaDCoWS2 clusters. Furthermore, the Euclid Wide Survey will largely overlap with the MaDCoWS2 footprint (Euclid Collaboration: Scaramella et al. 2022; Euclid Collaboration: Mellier et al. 2025).
In this paper, we present Euclid’s first view of a candidate cluster detected in both MaDCoWS2 and the Euclid preliminary cluster search: MOO2 J0337428386/EUCLQ1CL J033730.18283827.6. We nicknamed this candidate the ‘Puddle’ cluster due to the intriguing BCG complex revealed by Euclid. The left panel of Fig. 1 presents a tricolour image based on Euclid data; the right panel shows a zoomed view of the BCG complex. This system represents a test of the capabilities – and limitations – of Euclid for the discovery and follow-up of high-redshift galaxy clusters.
The paper is divided as follows: Sect. 2 presents the data used to discover and characterise the Puddle cluster. Section 3 presents the data analysis, focusing on the galaxy population and the central structure (i.e. the BCG complex) characterisation. In Sect. 4 we discuss our findings, which are summarised in Sect. 5. Throughout this paper, we assume a Planck Collaboration: Aghanim et al. (2020) cold dark matter (CDM) cosmology as implemented in astropy.cosmology: and . At , an angular distance of corresponds to a proper distance between 1.03 and 1.04 Mpc. All magnitudes are in the AB system. We assumed a Kroupa and Boily (2002) initial mass function.
2 Data
2.1 MaDCoWS2 catalogue
MaDCoWS2 is an optical and near-infrared catalogue that covers and contains 133 036 candidate galaxy clusters (Thongkham et al. 2024b, a). It is based on the bands from the DECam Legacy Survey (DECaLS; see Flaugher et al. 2015; Dey et al. 2019) and the bands from the CatWISE2020 catalogue (Eisenhardt et al. 2020; Marocco et al. 2021). The photometric redshift probability distribution functions (PDFs; see Brodwin et al. 2006) computed from these bands was input in PZWav. In MaDCoWS2, the Puddle cluster was detected as a overdensity with a signal-to-noise ratio (S/N) of 5.6.
2.2 Euclid Q1 data release
Euclid’s two instruments are the Visible Camera (VIS; Euclid Collaboration: Cropper et al. 2025), a single-band (denoted ) imaging camera covering the to Å range, and the Near-Infrared Spectrometer and Photometer (NISP; Euclid Collaboration: Jahnke et al. 2025). The NISP can perform imaging (, , and bands) and multiobject slitless spectroscopy (blue and red grisms).
The Euclid Consortium Quick Release 1 (hereafter the Q1 release, Euclid Quick Release Q1 2025; Euclid Collaboration: Mellier et al. 2025; Euclid Collaboration: Aussel et al. 2025) is composed primarily of Euclid images and photometric and spectroscopic catalogues (MER and SPE catalogues, see Euclid Collaboration: Aussel et al. 2025; Euclid Collaboration: Romelli et al. 2025; Euclid Collaboration: Copin et al. 2025) of the Euclid Deep Field North, Euclid Deep Field Fornax, and Euclid Deep Field South. The two last fields are also covered by the MaDCoWS2 catalogue. The Puddle cluster is located in the Euclid Deep Field Fornax.
A preliminary cluster search is also part of the Q1 release (Euclid Collaboration: Bhargava et al. 2025). This search focuses on detections. Euclid photometric redshifts are still being refined and currently have large uncertainties at . However, Euclid Collaboration: Bhargava et al. (2025) includes a sample of 15 unconfirmed clusters that meet the requirement of being detected both by PZWav and AMICO. The Puddle cluster is among this list. It was detected by PZWav as a overdensity with an S/N of 9.17 and by AMICO as a overdensity with an S/N of 16.39 and a richness of . AMICO and PZWAv compute S/N differently; hence AMICO produces a systematically higher S/N than PZWAv (see Fig. 4 of Euclid Collaboration: Bhargava et al. 2025). However, no redshift uncertainties are provided, and the scatter between photometric and spectroscopic redshifts is estimated for only. We therefore used the MaDCoWS2 detection as our primary reference for the cluster position and photometric redshift.
2.2.1 MER catalogue
The Q1 MER catalogue111available in Euclid Science Archive System, at https://eas.esac.esa.int/sas/ combines VIS, NISP, and several external data sets, including bands from DECam (Honscheid and DePoy 2008; Flaugher et al. 2015) on the 4 m Blanco Telescope, to create a common catalogue with consistently measured photometry. Euclid and external images are organised in calibrated tiles with overlaps and pixel scale (Euclid Collaboration: Romelli et al. 2025). Sources are detected in VIS and NISP, de-blended, and organised into a single source list which is then used to measure photometry in all bands.
To identify the presence of a red sequence, we used the and bands, which bracket the Å break at . While the band lies blueward of the Å break at these redshifts, its broadness (effective width of Å, with a mean wavelength of Å, see Rodrigo et al. 2012; Rodrigo and Solano 2020; Rodrigo et al. 2024) limits its usefulness for characterising galaxy populations. Specifically, we used the and fluxes obtained via a single Sérsic model-fitting, denoted by a ‘_sersic’ suffix in the MER catalogue.
2.2.2 MER images
The MER catalogue pipeline performs well for most galaxies, but tends to overcorrect or suppress low-surface brightness features like tidal tails (Euclid Collaboration: Aussel et al. 2025; Euclid Collaboration: Roster et al. 2025). Furthermore, the single Sérsic and circular aperture photometry currently available are poorly suited to irregular galaxies. Thus, rather than rely on the MER catalogue photometry of the BCG complex, we downloaded Euclid and DECam level 2 mosaics of the Puddle cluster, and trimmed them to a more manageable size of . We estimated the flux of the BCG with sep (Barbary 2018), using a custom aperture (see Sect. 3.3). The level 2 DECam mosaics in the Euclid Science Archive System are already resampled to the pixel scale of the Euclid images.
2.3 Spitzer data
The Spitzer Space Telescope (Werner et al. 2004) was a near- and mid-infrared observatory launched in August 2003 with three instruments on board: the Infrared Array Camera (IRAC, Fazio et al. 2004), the Multiband Infrared Photometer for Spitzer (MIPS; Rieke et al. 2004), and the Infrared Spectrograph (IRS, Houck et al. 2004). IRAC had four bands, usually called ‘channels’, centred on , , , and ; MIPS had three bands, centred on , , and . However, after Spitzer exhausted its liquid helium reserve in May 2009, only the two bluest bands of the IRAC camera could be used.
The Cosmic Dawn Survey (Euclid Collaboration: Moneti et al. 2022, see also Euclid Collaboration: Zalesky et al. 2025) is an IRAC survey covering the three Euclid Deep Fields and several calibration fields. Its image mosaics are based on both dedicated and archival data in the four IRAC channels, and correspond to about 11% of the telescope total mission time (Euclid Collaboration: Moneti et al. 2022). A source catalogue is available; however, the aperture sizes are inappropriate for the shape of the BCG. Furthermore, the BCG complex is blended with many neighbouring galaxies at and and is barely detected at and . We thus estimated the flux directly from the mosaics (see Sect. 3.3).
The coordinates of the brightest core in the BCG complex correspond ( separation) to a point source detected at in the Spitzer Wide-area InfraRed Extragalactic Survey (SWIRE; see Lonsdale et al. 2003, 2004). Given the relative isolation of this source (the closest MIPS source is away) and the FHWM of the MIPS point-spread function (PSF; Rieke et al. 2004), we adopted the SWIRE flux here.
2.4 Keck MOSFIRE spectroscopy
The Euclid Q1 release contains NISP slitless multiobject spectra taken with the red grism (Euclid Collaboration: Copin et al. 2025; Euclid Collaboration: Jahnke et al. 2025). However, the reduction of the grism data was not optimal in Q1; most of the identified issues are being addressed for the Euclid DR1. In the Puddle cluster field of view, an average of 55.6% of the pixels of the calibrated unidimensional spectra (excluding the edges) are flagged for suspicious behaviour.
Thus, to determine the Puddle cluster redshift, we obtained Keck MOSFIRE multiobject spectroscopy in the band. The observations were taken on the night of 24-25 December 2024, with a total on-target integration time of s. The weather was clear with an airmass between 1.5 and 1.7 during the observations, and the average seeing was about . We used s integrations on each node of a 2-point dither pattern with a separation between the nodes, and a single mask with 18 slits (usually ) for the whole observation. The targeted objects were selected to have magnitudes consistent with high-redshift galaxies in the MER catalogue and different colours. One of the slits was positioned across the BCG complex; its 1 and 2D spectra are presented in Sect. 3.2.
Spectroscopic data were reduced with the PypeIt data reduction pipeline version 1.17.1 (Prochaska et al. 2020), following the standard calibration and extraction process. The reduction process included bias subtraction, flat-field correction, and wavelength calibration employing arc lamp frames. Sky subtraction was performed using the PypeIt algorithm, which models the sky background on a slit-by-slit basis to account for variations across the field of view. PypeIt outputs reduced and co-added 2D spectra, from which we extract 1D spectra using a constant extraction width as a function of wavelength with an object-finding algorithm. The extraction width is calculated based on a PypeIt-based optimal aperture mask maximising the S/N. We examined the 1D and 2D spectra to locate emission lines.
3 Analysis
The detection of the Puddle cluster in MaDCoWS2 and Euclid (Euclid Collaboration: Bhargava et al. 2025) indicated the presence of a sizeable overdensity. MOSFIRE spectroscopy provided spectroscopic redshifts centred at for the BCG (see Sect. 3.2) and for a galaxy at away from the BCG complex, which may be part of the cluster infalling region. In this section, we explored the candidate cluster properties with the photometric and spectroscopic data sets. When necessary, we assumed that the redshifts of the cluster and the BCG are .
3.1 The galaxy population
Figure 2 presents the colour-magnitude diagram (CMD) of the Puddle cluster. The galaxies out to away (about Mpc) from the proto-BCG are shown as green points, and those between and are in black. The red sequence colour is 2.44 at (pink line), assuming Solar metallicity. It is computed with Bagpipes (Carnall et al. 2018) with a model consisting of a single burst of star formation at followed by passive evolution. Combining this model with the Mancone et al. (2010) luminosity function for the IRAC channel, we compute a characteristic magnitude () of 21.4 in band. A formation redshift of 3 generates similar results: a red sequence at and .
The central structure is detected as four galaxies by the MER pipeline. However, given the irregular shape and the presence of substantial diffuse emission, the MER photometry is unreliable for these four sources. Thus, we removed those four detections from our CMDs and performed our own measurement of the proto-BCG flux in Sect. 3.3. The cyan dot in Fig. 2 represents the position of the central structure in the CMD.
The red sequence of our CMD appears sparsely populated compared with clusters confirmed at (e.g. Papovich et al. 2010; Andreon et al. 2014; Webb et al. 2015b; Strazzullo et al. 2016; Willis et al. 2020). We see in particular a lack of red galaxies at , which is unusual among the known high-redshift clusters: massive galaxies are more susceptible to internal quenching (e.g. Peng et al. 2010b) and thus tend to be redder than less massive ones even at (e.g. Kawinwanichakij et al. 2017).
| Decomposition | Line | FWHM | AIC | ||
| One Gaussian | H | 1.82 | 3263 | ||
| Two Gaussians | H | 1.61 | 3190 | ||
| H | |||||
| Three Gaussians | [N ii] | 1.62 | 3194 | ||
| H | |||||
| [N ii] |
3.1.1 Background subtraction and completeness correction
To assess the importance of the galaxy overdensity associated with the Puddle cluster and its location in the CMD diagram, we estimated the contribution of field galaxies and subtracted it from our CMD. To measure this ‘background’, we randomly selected 100 non-overlapping circular regions () in the Fornax field. We computed the average density of galaxies in these fields as a function of magnitude and colour and we used only the galaxies that are at least -detected in both and bands. For each magnitude and colour, we then subtracted this average density from the galaxy density in the central region of the Puddle cluster. The background subtraction uncertainties were determined from the standard deviation of the galaxy densities in the random fields.
We also used the 100 random fields described above to determine completeness corrections. Our computation followed the same principle than the completeness correction described in the Appendix A.1. of van der Burg et al. (2020). Figure 3 presents a visual summary of the process. We fitted an exponential to the highly complete part of the distribution in each band empirically determined as and . It should be noted that since the -band completeness is the most critical one, we have been especially conservative in our selection of the regime which we consider highly complete. We then computed a completeness factor for magnitudes greater than 23 () or 23.5 () by taking the ratio of the observed galaxy count to the count predicted by the best fit. We assumed a completeness factor of one for and . The total 50% completeness, shown on Fig. 2 as a purple line, is estimated by multiplying the and -band completeness factor for a given set of and magnitudes. This multiplication relies on the assumption that there is no correlation between detections in the and bands. Since that may not be true, we limited the rest of our analysis to , a regime where the completeness correction does not fall below 34% and is entirely determined by the -band completeness (i.e. the completeness correction is entirely determined by colour). The completeness uncertainties were propagated from the uncertainties on the best fit parameters. For example, the correction for a -band completeness of 0.5 is .
Figure 4 presents the impact of these corrections on the galaxy population of the Puddle cluster candidate. The bottom left panel shows the background-subtracted, completeness-corrected distribution. The 50% -band completeness limit is shown as a cyan line. To avoid large (and uncertain) completeness corrections, we limit the CMD in the bottom left panel to ( is equivalent to on the red sequence and to a 76% complete band). We limited the colour to , as the background-subtracted galaxy count is consistent with zero for redder colours. The other panels show the galaxy count before and after the application of the completeness correction as a function of (top left) and colour (bottom right). For , the excess count within of the BCG is galaxies above the field level. It increases to galaxies after the application of the completeness correction. The red sequence location is indicated on the top left and bottom panels. The count uncertainties are based on the propagation of the density subtraction and the completeness uncertainties when relevant.
As noted above, there are galaxies ( after the completeness correction) in excess compared to the field level in the red sequence and very few luminous red galaxies outside of the BCG complex ( galaxies with ). The bulk of the overdensity is about 1.5 magnitudes bluer than the red sequence population, and fainter than . With Bagpipes, we computed that corresponds to a Myr old stellar population, assuming Solar metallicity and no attenuation. While this blue colour suggests widespread star formation, we did not attempt to determine the cluster specific SFR. The lack of bright red galaxies and the imprecise photometric redshifts of Euclid at , are probably the two main factors explaining the difference between our galaxy counts and the Euclid Collaboration: Bhargava et al. (2025) richness estimate.
3.2 Spectroscopy
We detected continuum emission of varying robustness in 14 of the 18 MOSFIRE slitlets. We also identified unambiguous isolated emission lines, most likely H , in three slitlets: one slitlet targeting the BCG complex, one slitlet targeting a source approximately from the MaDCoWS2 centroid, and another slitlet far from the MaDCoWS2 centroid that has two serendipitous emission lines. In addition, faint, tentative emission lines were detected in three other slitlets. Based on the one of these faint tentative line, we estimated a line detection threshold of at the central wavelengths of the -band grism. This corresponds to a star-formation rate of , assuming a standard conversion factor between star-formation rate and H line luminosity (Hao et al. 2011). At m, the wavelength of the emission line detected from the BCG complex, the atmosphere absorbs of the flux, implying a rough star-formation rate detection threshold of 8. The presence of faint traces and tentative emission lines suggests that the exposure time calculations were perhaps too optimistic. We also note that the outcome of our observation is comparable to Stanford et al. (2012) attempt to confirm IDCS J1426.5+3508, a cluster, with spectra from the Low-Resolution Imaging Spectrograph on Keck I. On 16 targets, they confirmed two members. However, the main challenge in our data reduction was the subtraction of the strong near-infrared sky lines (e.g. Webb et al. 2015b; Balogh et al. 2017; Belli et al. 2019), while Stanford et al. (2012) main limitation was the faintness of quiescent galaxies in the optical (which corresponds to rest-frame ultraviolet at ).
Figure 5 presents a portion of the spectrum of BCG complex, centred on the emission line. As shown by the top panel, the emission originates solely from the proto-BCG central nucleus. While the slit was designed to be centred on the main core and to include the southwestern nucleus (see the right panel of Fig. 1) we do not see any clear secondary trace. It is possible that the southwestern nucleus may possess only a very faint continuum and no significant emission lines; we note also that the extensive diffuse emission in the BCG complex may have been subtracted as sky emission, resulting in a lower signal in the vicinity of the southwestern nucleus. Alternatively, if a small misalignment occurred during observations, the second nucleus may have fallen outside of the slit during some of the dithers. The spectral resolution varies across the spectrum, but is on average Å pixel-1.
The emission region in Fig. 5 is broad, spanning roughly Å and presents a complex structure. Such a broad range suggests three possibilities: a single H line from an AGN; at least two sources of H emission along the same line-of-sight; or H emission blended with the [N ii] doublet. We examined these possibilities by modelling the spectrum as one, two, and three Gaussians, weighting the contribution of each pixel by the inverse of its variance. We computed the reduced and Akaike information criterion (AIC, e.g. Akaike 1974; Liddle 2007) of each fit. The results are presented in Table 1. The middle and lowest panels of Fig. 5 show the one-, two-, and three-Gaussian decompositions in blue, pink, and purple, respectively. Due to the limited S/N of the spectrum, we did not attempt to subdivide the H emission into broad and narrow components.
The simplest case, the single Gaussian model, yields the highest reduced and AIC. Both measures strongly suggest that the two- or three-Gaussians models are better representations of the spectrum. The components of the three-Gaussian model are not treated independently: We fixed the amplitude ratio of the [N ii] doublet as , in agreement with theoretical (e.g. Galavís et al. 1997; Storey and Zeippen 2000; Bon et al. 2025) and observational (Dočinović et al. 2023) determinations, and forced its members to have the same width. We forced all three emission lines to have the same redshifts. We also tested a fit in which we relaxed the last requirement and allowed the H emission to lie at a different redshift than the [N ii] doublet. We found very similar redshifts () and the FWHM of each line was not significantly changed compared to the previous fit. We discussed the interpretation of the various fits in Sect. 4.2.1.
3.3 The BCG complex
The right panel of Fig. 1 presents a tricolour image (, , and ) of the BCG complex of the Puddle cluster, with the MOSFIRE slit and the custom aperture used for the photometry overlaid (see Sect. 3.3.1). This system has a complex morphology, with diffuse emission embedding at least six cores and extending in a kind of tail towards the southwest. There is a seventh core within the tail emission. The projected distance between cores is kpc, and in the case of the six central cores kpc. Along the axis of the diffuse tail, the system is kpc across.
3.3.1 Euclid and DECam photometry
To estimate the magnitude of the BCG complex, our first step was to determine a suitable aperture to measure the system photometry. To do so, we used sep.extract to compute the segmentation map in the NISP band and set our detection threshold at with a minimum area of 30 pixels. We then assigned a value of one to the pixels corresponding to the BCG complex detections and removed all other detections in our segmentation map. The resulting detection was very irregularly shaped, so we used a Gaussian kernel with 3 pixels to smooth the detection shape, retaining only the pixels with a value of 0.25 or more. We then used the resulting shape (shown as white contours in the right panel of Fig. 1) to measure the photometry in all Euclid and DECam bands.
To estimate the photometric uncertainties, we adopted a statistical approach: We found all detections in the image of interest, with a detection threshold at and a minimum area of 5 pixels. We then randomly displaced the footprint of the central merger until we found a position with no overlapping detections. We measured the sky flux in that position and repeated the process for 100 random empty sky patches. We took the standard deviation of the sky fluxes as our sky uncertainties, which we then added in quadrature to the Poisson noise of the aperture photometry.
We did not compute aperture corrections for DECam. As long as the defined aperture contains most of the flux in every band, the aperture corrections should be negligible. This is the case for all DECam bands. In 3 segmentation maps, the BCG complex, if detected, is entirely enclosed by our custom aperture.
3.3.2 IRAC photometry and morphology modelling
The pixel size of the IRAC mosaics is , which is six times larger than the the pixel size of Euclid VIS and NISP images. To measure the flux of the marginal detections in the IRAC and channels, we resampled our convolved segmentation map to the correct pixel size (setting the pixels with fractional values to unity) and performed the flux and uncertainty measurements as described above.
However, at and the BCG complex is heavily blended with neighbouring galaxies. We used Galfit modelling (Peng et al. 2002, 2010a) to disentangle the structure from neighbouring galaxy fluxes. Using the high resolution of the NISP data (Fazio et al. 2004; Euclid Collaboration: Jahnke et al. 2025), we constructed initial models in band. We first modelled the BCG complex with seven Sérsic profiles for the structure nuclei, and two additional Sérsic profiles for the diffuse emission. Because no PSF estimation is provided with Q1 frames, we approximated Euclid PSFs using a neighbouring unsaturated star. The result is shown in Fig. 6. We then made a separate model of the contaminants before merging the two models together.
Our IRAC and models are directly based on the combined -band models, with a few changes. Unlike the Euclid data, IRAC mosaics have uncertainty maps, which we added into the Galfit settings. We also deleted one of the two diffuse Sérsic profiles, as we could not make the merged models converge otherwise. Likewise, we removed some of the faintest interlopers, as they were too faint to be detected in the IRAC images. For our first pass models, we only let the magnitudes change, but for our final models we allowed positional changes as well to account for shifts induced by PSF miscentring; the effective radii, Sérsic indexes, axis ratios, and position angles were fixed to initial values.
We tested two different PSFs: an empirical PSF based on the same star used to estimate the -band PSF, and a stacked theoretical point response function (PRF). The stacked PRF was made by rotating the model PRF of the cryogenic mission222see https://irsa.ipac.caltech.edu/data/SPITZER/docs/irac/calibrationfiles/psfprf/ to the position angles of all the input images contributing to the mosaic at the location of the BCG. The rotated PRFs were weighted by the square root of the exposure times and stacked. The empirical PSF performed better than the stacked PRF, which was too concentrated.
We measured the flux of the BCG complex in the IRAC and channels directly from the best Galfit models. To ensure consistency with the aperture photometry (see Sect. 3.3.1), we computed an aperture correction by dividing the aperture flux of the band by the Galfit flux in the same band. We found an aperture correction of .
3.3.3 Spectral energy distribution fitting
| Instrument | Band | Flux |
| DECam | ||
| DECam | ||
| DECam | ||
| DECam | ||
| VIS | ||
| NISP | ||
| NISP | ||
| NISP | ||
| IRAC | ||
| IRAC | ||
| IRAC | ||
| IRAC | ||
| MIPS |
1 After multiplication by an aperture correction factor of . See the main text for more details.
Table 2 presents the derived DECam, Euclid, and IRAC photometry, which we used to perform spectral energy distribution (SED) modelling of the entire BCG complex. We did not attempt to model the SED of individual cores: The signal-to-noise of the and DECam fluxes was not sufficient to perform a Galfit decomposition.
We adopted most of the Bagpipes (Carnall et al. 2018) standard basic assumptions (models from Bruzual and Charlot 2003, with a Kroupa and Boily 2002 initial mass function, uniform priors), and added a dust component that follows the Calzetti et al. (2000) attenuation law. We replaced the Bagpipes default cosmology with the Planck Collaboration: Aghanim et al. (2020) cosmological parameters. We adopted a parametric approach, testing a delayed -model and a model in which all the stars formed at the same time (i.e. a burst). The burst yields a best fit SED with a smaller reduced than the delayed -model (3.34 compared to 3.75). Furthermore the characteristic time of the delayed -model is extremely short, Myr. The free parameters, priors, and results of the bursty model are presented in Table 3. The results are expressed as the median of the parameter probability distribution, and their uncertainties were estimated using the 16 and 84th percentiles. The resulting SED is shown in Fig. 7. Note that Bagpipes fits the total stellar mass formed. The current stellar mass is , and is also included in Table 3.
We also tested more complicated models: first two bursts and then a burst with a delayed -model, assuming Solar metallicities to reduce the number of free parameters. In both cases, the results were very poorly constrained.
In all these fits, we assumed that the central AGN contribution to the flux is minimal. To assess the impact of the AGN emission on the fit, we used the AGN module in Bagpipes (Carnall et al. 2023), which implements a Vanden Berk et al. (2001) quasar model. The model consists of two power laws describing the continuum blueward and redward of 5100 Å in the rest frame, normalised by the flux of the AGN at 5100 Å and by the flux and width of the H emission line. We tested different variations of this model, assuming that the AGN is responsible for 10% of the flux at 5100 Å (which we estimate using the most likely SED of the fit without an AGN) and that the flux of the H emission line corresponds to 10% of the band flux. We obtained large reduced values. We then let the 5100 Å normalisation vary and found that the most likely AGN contribution to the flux at 5100 Å is negligible (%) and that the other parameters of the fit corresponded to the values listed in Table 3. We thus concluded that the AGN does not contribute significantly to the SED of the BCG complex.
| Parameter | Prior1 | Results |
| Age (Gyr) | [0.001,3.74]2 | |
| Mass formed [] | [10,13] | |
| Av (mag) | [0,2] | |
| Metallicity ( | [0,2.5]3 | |
| Stellar mass []4 | - |
1 All priors are uniform.
2 The code automatically limited the prior to the age of the Universe at the considered redshift.
3 We also tested models with fixed metallicities. See the explanation in the main text.
4 Bagpipes does not fit directly the current stellar mass, but rather the total stellar mass formed during the star-formation history. We included the current stellar mass for indicative purposes only.
4 Discussion
4.1 Cluster red fraction
The Puddle is a candidate cluster detected in two different surveys, MaDCoWS2 and Euclid (Thongkham et al. 2024a; Euclid Collaboration: Bhargava et al. 2025). It still corresponds to a significant galaxy overdensity after subtraction of the mean galaxy density in the Euclid Deep Fornax Field (see Sect. 3.1). While a single redshift is not enough to spectroscopically confirm a cluster (e.g. Eisenhardt et al. 2008; Noirot et al. 2018), the facts enumerated above suggest that the Puddle cluster is a genuine high-redshift cluster (see also the characterisation of similar unconfirmed clusters by Valtchanov et al. 2004 and Cooke et al. 2016).
As noted in Sect. 3.1, the Puddle cluster red sequence is relatively sparse compared to other clusters and dominated by faint red members. This situation would make the detection of the Puddle cluster challenging with the red-sequence method (e.g. Gladders and Yee 2000, 2005; Donahue et al. 2002), one of the commonly used algorithms for optical/near-infrared cluster searches (e.g. the C4 clustering algorithm, the Spitzer Adaptation of the Red-sequence Survey and the red-sequence Matched-filter Probabilistic Percolation; Miller et al. 2005; Wilson et al. 2006, 2009; Rykoff et al. 2014).
While Fig. 2 does not show a clear division between the red sequence and the blue cloud, in the colour histograms shown in Fig. 4 bottom right panel, there is a local minimum at . We thus adopted , which corresponds to a stellar age of 1.08 Gyr (Solar metallicity, no attenuation), as our limit between the red sequence and the blue cloud. With this assumption, we found, within a radius centred on the BCG complex, red galaxies () without completeness correction, and red galaxies with it. In both cases, the count of blue galaxies was . These counts correspond to red fractions of % and % respectively. The uncertainties were based on the propagation of the uncertainties of the background subtraction and the completeness correction (see Sect. 3.1.1).
This would make the red fraction of the Puddle cluster consistent with the measurements of Quadri et al. (2012) and Nantais et al. (2016) in clusters, but significantly less quenched than the clusters of Newman et al. (2014), Cooke et al. (2016), and Strazzullo et al. (2019). However, the comparison with other works is made challenging by the different stellar mass, magnitude, and radius cuts applied by each author, and by the dust obscuration, which can be mistaken for quiescence (e.g. Woo et al. 2013; Hatch et al. 2016).
In particular, our own magnitude cut does not correspond to a uniform stellar mass cut. At similar -band magnitudes, blue galaxies will have a lower stellar mass than red members. Assuming a simple stellar population model with no reddening, we computed approximate stellar mass as a function of colour and magnitude and found no overlap between the mass range probed by the reddest and the bluest colour in our CMD. To enable a uniform stellar mass cut, we would need deeper -band data than those currently available.
4.2 BCG formation scenarios
The morphology of the BCG complex is consistent with a complex multiobject merger. While line-of sight superpositions remain possible, the six central cores are embedded in diffuse emission, which suggests that they are interacting. The southwestern core may or may not be part of the merger. The extension of the diffuse emission to the southwest is reminiscent of a tidal tail (see e.g. Webb et al. 2015b). We thus inferred that the Puddle cluster is in the process of forming its BCG through a multiobject merger. In this section, we explored the properties of this merger and how it compares with SPT234956 (Miller et al. 2018) and SpARCS104922.6+564032.5 (Webb et al. 2015b). We noted two other instances of multiobject mergers interpreted as assembling BCGs in the literature: the Spiderweb protocluster at (Miley et al. 2006) and an unnamed overdensity in the Cosmic Evolution Early Release Science of JWST (Coogan et al. 2023).
4.2.1 Star formation and AGN activity
Most of the proto-BCGs in the literature exhibit high levels of star formation, obscured or not (Seymour et al. 2012; Webb et al. 2015b; Miller et al. 2018; Coogan et al. 2023). In the Puddle cluster, the SED model that best fits the photometry of the BCG complex is an instantaneous burst of star formation Myr ago, with no more recent star formation. We note however that parametric SED modelling should be taken with caution (e.g. Carnall et al. 2019; Leja et al. 2019; Haskell et al. 2024; Mosleh et al. 2025). Thus, we interpret these results as evidence for a burst of star formation about Myr ago, but do not rule out the possibility of more recent, residual star formation. In fact, given the young age of the dominant stellar population, ongoing unobscured star formation in the Puddle cluster proto-BCG is plausible. Obscured star formation is also possible: All star-formation histories returned large dust attenuations, regardless of the metallicity content. Without coverage in the far infrared, it is difficult to assess its importance.
The spectrum of the brightest nucleus does not offer any clues about the current SFR since it is dominated by AGN emission. If we model the spectrum as two Gaussians (see Sect. 3.2), the broadest emission line FWHM is (i.e. ). This is too large to be associated with the merging activity, since the line width is similar to the total velocity dispersion of an entire (proto)cluster (e.g. Kuiper et al. 2011; Ruel et al. 2014).
The results of the three-Gaussian modelling also suggest that the spectrum is AGN-dominated. Using this model, we computed approximate equivalent-width ratios of and . Large values like these are unlikely to be associated with star formation alone, but consistent with the ionisation provided by an AGN (e.g. Bian et al. 2018; Oh et al. 2019, 2022; Garg et al. 2022; Dors et al. 2023; Euclid Collaboration: Scharré et al. 2024; Zhou et al. 2025).
The stellar mass computed by Bagpipes is . This value does not include AGN contamination and assumes a single stellar population, about 300 Myr old. An AGN can sometimes result in an overestimation of the stellar mass (e.g. Ciesla et al. 2015; Salim et al. 2016). However, the presence of an older population would instead decrease the mass-to-light ratio (e.g. Mancone and Gonzalez 2012; Conroy 2013). Bagpipes stellar mass should thus be treated as an estimate.
4.2.2 Comparison with SPT234956 simulations
Lying at , SPT234956 is a protocluster with 14 massive gas-rich galaxies all within kpc of projected distance from each other (Miller et al. 2018). According to the simulations of Rennehan et al. (2020, see also updated versions), SPT234956’s 14 central galaxies should merge into a single BCG in Myr from the time of observation. They found that the star-formation rate peaks very early in their simulation, before decreasing exponentially with a characteristic time of Myr. Remus et al. (2023) comparison with the Magneticum cosmological simulations also predicts a rapid collapse.
The proto-BCG of the Puddle cluster has not yet coalesced into a single galaxy. However, the age and burst-like star-formation history of the unobscured stellar population suggest that we could be witnessing a more advanced stage of an SPT2349-like merger. The more compact morphology of the Puddle cluster’s proto-BCG (see Sect. 3.3) also supports this hypothesis. Likewise, the stellar mass computed by Bagpipes, , is larger than the initial stellar mass of the Rennehan et al. (2020) simulations () and smaller than the predicted mass in Myr ().
Rennehan et al. (2020) also made predictions about the probability of observing massive collapse events like SPT234956 (or the Puddle cluster) at different redshifts. While they observed events at all redshifts, most of them occured before . Yet, observationally, SPT234956 is the highest-redshift merging proto-BCG in the literature (Miller et al. 2018). All the others lie at redshifts closer to the Puddle cluster – between and (Miley et al. 2006; Webb et al. 2015b; Coogan et al. 2023). A sample of five proto-BCGs is not enough to draw statistically significant conclusions. However, the fact that most of them lie at suggests either that most collapses occur later and/or on longer timescales than the Rennehan et al. (2020) predictions, or that selection effects decrease the likelihood of detecting merging BCGs at .
4.2.3 Comparison with SpARCS104922.6+564032.5
While SPT234956 is the only proto-BCG with dedicated simulations, the SpARCS104922.6+564032.5 system is more directly comparable to the Puddle cluster in terms of redshift and physical scale. At , SpARCS104922.6+564032.5’s star-forming core is dominated by a kpc-long tidal tail, with small clumps of various colours (Webb et al. 2015b). It is unclear if the tidal tail is directly associated with the BCG, or with the pair of smaller merging cluster members located to its north. Subsequent studies have shown that the BCG hosts a faint radio-loud AGN (Trudeau et al. 2019) and that the bulk of the star formation occurs at kpc to the southeast of the BCG (Barfety et al. 2022). This corresponds to the centre of the tidal tail, where the MIPS emission of the system is located. The discovery of X-ray emission centred on the star-forming region suggests that this system might be powered by a displaced cooling flow rather than a merger (Hlavacek-Larrondo et al. 2020).
The presence of a cooling flow in the Puddle cluster cannot be determined without X-ray or SZ observations. However, the Puddle cluster displays none of the clues that suggested the presence of a cooling flow as in SpARCS104922.6+564032.5. Unlike that cluster, the BCG of the Puddle cluster is clearly merging. Furthermore, the centroid of the MIPS emission is almost directly located onto the brightest of the merging galaxies. The proto-BCG emission lines are most likely powered by an AGN (see Sect. 4.2.1) and are inconsistent with star formation as the primary ionising mechanism. The H line is also too broad to be produced by a filamentary nebula similar to those associated with local cooling flows (e.g. McDonald et al. 2010; Hamer et al. 2016; Gaspari et al. 2018; Gendron-Marsolais et al. 2018; Olivares et al. 2019, 2022; Mohapatra et al. 2022). Only the extreme cooling flow in the Phoenix cluster – which is associated with a large amount of unobscured star formation (McDonald et al. 2012, 2013) – produces emission lines with FWHM around (Reefe et al. 2025).
4.3 Finding other assembling BCGs
The discovery of an assembling proto-BCG in one of the two Euclid Deep Fields covered by MaDCoWS2 suggests that these objects might be common – a hypothesis that is further supported by the presence of similar merging BCGs in the literature. Yet it is difficult to assess the prevalence of these systems because high-resolution near-infrared (e.g. this study; see also Kuiper et al. 2011; Coogan et al. 2023) or (sub)millimetre (Miller et al. 2018, see also Rotermund et al. 2021) observations are necessary to reveal the complex structures of proto-BCGs.
The large areas of the sky that will be observed by the Euclid, Nancy Grace Roman, and Vera C. Rubin should allow more systematic detections of merging proto-BCGs. In particular, the Euclid Wide Survey will cover approximately deg2 at the same depth as the Q1 data release (Euclid Collaboration: Scaramella et al. 2022; Euclid Collaboration: Mellier et al. 2025; Euclid Collaboration: Aussel et al. 2025). Thus, assuming that our single detection in the overlap between Euclid Q1 Data Release and MaDCoWS2 (deg2) is representative of the density of proto-BCGs on the sky, we would expect to find about 400 proto-BCGs in the completed Euclid Wide Survey.
5 Conclusions
This work presents the characterisation of MOO2 J0337428386/EUCLQ1CL J033730.18283827.6 (nicknamed the ‘Puddle’ cluster), a galaxy cluster candidate with a merging proto-BCG.
The Puddle cluster is a candidate cluster in MaDCoWS2 catalogue and is detected as a candidate cluster by both Euclid cluster-finding algorithms. Compared to field level, there is an excess of galaxies brighter than within a radius around the proto-BCG. With a completeness correction, this number rises to galaxies. This excess of galaxies strongly suggests that this candidate cluster is a genuine high-redshift overdensity, even though only the BCG is spectroscopically confirmed at with H emission.
The Puddle cluster central does not appear to contain luminous () red galaxies aside from the BCG complex. We estimate the fraction of red galaxies ( and ) as % without the completeness correction and as % with it. It should be noted however that does not correspond to a uniform stellar mass – deeper -band data would be needed to compute stellar-mass based red fractions.
The proto-BCG appears to be a complex merger at , including 6 to 7 galaxies. Along its longest axis, the BCG system measures about kpc, including its southwestern tail of diffuse emission. The emission spectrum associated with this system suggests that the brightest of the merging cores hosts an AGN.
The results of SED modelling of the forming BCG indicate that the dominant stellar population is young (Gyr) and consistent with a short burst of star formation. The central AGN contribution to the SED is negligible. However, the most likely SEDs show substantial dust attenuation. Thus, more data are needed to determine the amount, if any, of obscured star-formation activity. The proto-BCG has a stellar mass of .
The morphology, size, and stellar mass of the proto-BCG, as well as the timescale of the star formation burst, suggest that the Puddle cluster could be a SPT234956-like merger, caught at a more advanced stage.
Acknowledgements.
We thank C.C. Steidel for providing an archival sensitivity function for MOSFIRE. A.T., A.H.G., S.A.S., S.T., D.S., P.R.M.E and B.M. acknowledge the support of NASA ROSES Grant 12-EUCLID11-0004. The work of D.S. and P.R.M.E. was carried out at the Jet Propulsion Laboratory, California Institute of Technology, under a contract with the National Aeronautics and Space Administration (80NM0018D0004). B.M. acknowledges funding support from the Jet Propulsion Lab (JPL) Euclid project. This work has made use of the Euclid Quick Release Q1 data from the Euclid mission of the European Space Agency (ESA), 2025, https://doi.org/10.57780/esa-2853f3b. The Euclid Consortium acknowledges the European Space Agency and a number of agencies and institutes that have supported the development of Euclid, in particular the Agenzia Spaziale Italiana, the Austrian Forschungsförderungsgesellschaft funded through BMIMI, the Belgian Science Policy, the Canadian Euclid Consortium, the Deutsches Zentrum für Luft- und Raumfahrt, the DTU Space and the Niels Bohr Institute in Denmark, the French Centre National d’Etudes Spatiales, the Fundação para a Ciência e a Tecnologia, the Hungarian Academy of Sciences, the Ministerio de Ciencia, Innovación y Universidades, the National Aeronautics and Space Administration, the National Astronomical Observatory of Japan, the Netherlandse Onderzoekschool Voor Astronomie, the Norwegian Space Agency, the Research Council of Finland, the Romanian Space Agency, the Swiss Space Office (SSO) at the State Secretariat for Education, Research, and Innovation (SERI), and the United Kingdom Space Agency. A complete and detailed list is available on the Euclid web site (www.euclid-ec.org/consortium/community/). This project used data obtained with the Dark Energy Camera (DECam), which was constructed by the Dark Energy Survey (DES) collaboration. Funding for the DES Projects has been provided by the U.S. Department of Energy, the U.S. National Science Foundation, the Ministry of Science and Education of Spain, the Science and Technology Facilities Council of the United Kingdom, the Higher Education Funding Council for England, the National Center for Supercomputing Applications at the University of Illinois at Urbana-Champaign, the Kavli Institute of Cosmological Physics at the University of Chicago, Center for Cosmology and Astro-Particle Physics at the Ohio State University, the Mitchell Institute for Fundamental Physics and Astronomy at Texas AM University, Financiadora de Estudos e Projetos, Fundao Carlos Chagas Filho de Amparo, Financiadora de Estudos e Projetos, Fundao Carlos Chagas Filho de Amparo Pesquisa do Estado do Rio de Janeiro, Conselho Nacional de Desenvolvimento Cientfico e Tecnolgico and the Ministrio da Cincia, Tecnologia e Inovao, the Deutsche Forschungsgemeinschaft and the Collaborating Institutions in the Dark Energy Survey. The Collaborating Institutions are Argonne National Laboratory, the University of California at Santa Cruz, the University of Cambridge, Centro de Investigaciones Enrgeticas, Medioambientales y TecnolgicasMadrid, the University of Chicago, University College London, the DES-Brazil Consortium, the University of Edinburgh, the Eidgenssische Technische Hochschule (ETH) Zrich, Fermi National Accelerator Laboratory, the University of Illinois at Urbana-Champaign, the Institut de Cincies de l’Espai (IEEC/CSIC), the Institut de Fsica d’Altes Energies, Lawrence Berkeley National Laboratory, the Ludwig-Maximilians Universitt Mnchen and the associated Excellence Cluster Universe, the University of Michigan, the National Optical Astronomy Observatory, the University of Nottingham, the Ohio State University, the University of Pennsylvania, the University of Portsmouth, SLAC National Accelerator Laboratory, Stanford University, the University of Sussex, and Texas AM University. This work is based in part on observations made with the Spitzer Space Telescope, which was operated by the Jet Propulsion Laboratory, California Institute of Technology under a contract with NASA. This research has made use of the NASA/IPAC Infrared Science Archive, which is funded by the National Aeronautics and Space Administration and operated by the California Institute of Technology. Some of the data presented herein were obtained at Keck Observatory, which is a private 501(c)3 non-profit organization operated as a scientific partnership among the California Institute of Technology, the University of California, and the National Aeronautics and Space Administration. The Observatory was made possible by the generous financial support of the W. M. Keck Foundation. This research has made use of the SVO Filter Profile Service ‘Carlos Rodrigo’, funded by MCIN/AEI/10.13039/501100011033/ through grant PID2023-146210NB-I00.References
- A New Look at the Statistical Model Identification. ITAC 19, pp. 716–723. Cited by: §3.2.
- Star Formation and AGN Activity in Galaxy Clusters from z=1-2: a Multi-Wavelength Analysis Featuring Herschel/PACS. ApJ 825, pp. 72. External Links: ISSN 0004-637X, Link, Document Cited by: §1.
- Cosmological Parameters from Observations of Galaxy Clusters. ARA&A 49, pp. 409–470. External Links: ISSN 0066-4146, Link, Document Cited by: §1.
- A systematic search for galaxy proto-cluster cores at z ∼ 2. MNRAS 496 (3), pp. 3169–3181. External Links: ISSN 0035-8711, Link, Document Cited by: §1.
- JKCS 041: a Coma cluster progenitor at z = 1.803. A&A 565, pp. A120. External Links: ISSN 0004-6361, Link, Document Cited by: §3.1.
- The K-band Hubble diagram for the brightest cluster galaxies: a test of hierarchical galaxy formation models. MNRAS 297, pp. 427–434. External Links: ISSN 0035-8711, Link, Document Cited by: §1.
- The Bimodal Galaxy Color Distribution: Dependence on Luminosity and Environment. ApJ 615, pp. L101–L104. External Links: ISSN 0004-637X, Link, Document Cited by: §1.
- Gemini Observations of Galaxies in Rich Early Environments (GOGREEN) I: survey description. MNRAS 470, pp. 4168–4185. External Links: ISSN 0035-8711, Link, Document Cited by: §3.2.
- Evidence for a change in the dominant satellite galaxy quenching mechanism at z = 1. MNRAS 456, pp. 4364–4376. External Links: ISSN 0035-8711, Link, Document Cited by: §1.
- The colour of galaxies in distant groups. MNRAS 398, pp. 754–768. External Links: ISSN 0035-8711, Link, Document Cited by: §1.
- SEP: Source Extraction and Photometry. Astrophysics Source Code Library. Note: Astrophysics Source Code Library, record ascl:1811.004 External Links: 1811.004 Cited by: §2.2.2.
- An Assessment of the In Situ Growth of the Intracluster Light in the High-redshift Galaxy Cluster SpARCS1049+56. ApJ 930, pp. 25. External Links: ISSN 0004-637X, Link, Document Cited by: §4.2.3.
- AMICO: optimized detection of galaxy clusters in photometric surveys. MNRAS 473, pp. 5221–5236. External Links: ISSN 0035-8711, Link, Document Cited by: §1.
- MOSFIRE Spectroscopy of Quiescent Galaxies at 1.5 < z < 2.5. II. Star Formation Histories and Galaxy Quenching. ApJ 874, pp. 17. External Links: ISSN 0004-637X, Link, Document Cited by: §3.2.
- “Direct” Gas-phase Metallicity in Local Analogs of High-redshift Galaxies: Empirical Metallicity Calibrations for High-redshift Star-forming Galaxies. ApJ 859 (2), pp. 175 (en). External Links: ISSN 0004-637X, 1538-4357, Link, Document Cited by: §4.2.1.
- Galaxy Clusters Discovered via the Sunyaev-Zel’dovich Effect in the 2500-Square-Degree SPT-SZ Survey. ApJS 216, pp. 27. External Links: ISSN 0067-0049, Link, Document Cited by: §1.
- Ratios of forbidden [OIII] λλ4959,5007 and [NII] λλ6548,6583 lines in nearby narrow emission line galaxies. FrASS 12 (English). External Links: ISSN 2296-987X, Link, Document Cited by: §3.2.
- Red but not dead: unveiling the star-forming far-infrared spectral energy distribution of SpARCS brightest cluster galaxies at 0 < z < 1.8. MNRAS 469, pp. 1259–1281. External Links: ISSN 0035-8711, Link, Document Cited by: §1.
- How filaments of galaxies are woven into the cosmic web. Nat 380, pp. 603–606. External Links: ISSN 0028-0836, Link, Document Cited by: §1.
- Photometric Redshifts in the IRAC Shallow Survey. ApJ 651, pp. 791–803. External Links: ISSN 0004-637X, Link, Document Cited by: §1, §2.1.
- The Era of Star Formation in Galaxy Clusters. ApJ 779 (2), pp. 138 (en). External Links: Link, Document Cited by: §1.
- Stellar population synthesis at the resolution of 2003. MNRAS 344, pp. 1000–1028. External Links: ISSN 0035-8711, Link, Document Cited by: §3.3.3.
- The SRG/eROSITA All-Sky Survey. The first catalog of galaxy clusters and groups in the Western Galactic Hemisphere. A&A 685, pp. A106. External Links: ISSN 0004-6361, Link, Document Cited by: §1.
- The eROSITA Final Equatorial-Depth Survey (eFEDS). Galaxy clusters and groups in disguise. A&A 661, pp. A10. External Links: ISSN 0004-6361, Link, Document Cited by: §1.
- Growth of brightest cluster galaxies via mergers since z=1. MNRAS 434, pp. 2856–2865. External Links: ISSN 0035-8711, Link, Document Cited by: §1.
- The radio properties of cD galaxies in Abell clusters. I - an X-ray selected sample. AJ 99, pp. 14–30. External Links: ISSN 0004-6256, Link, Document Cited by: §1.
- The Dust Content and Opacity of Actively Star-forming Galaxies. ApJ 533, pp. 682–695. External Links: ISSN 0004-637X, Link, Document Cited by: §3.3.3.
- How to measure galaxy star-formation histories I: Parametric models. ApJ 873 (1), pp. 44. Note: Comment: 20 pages, 12 figures, ApJ accepted External Links: ISSN 1538-4357, Link, Document Cited by: §4.2.1.
- Inferring the star formation histories of massive quiescent galaxies with BAGPIPES: evidence for multiple quenching mechanisms. MNRAS 480, pp. 4379–4401. External Links: ISSN 0035-8711, Link, Document Cited by: §3.1, §3.3.3.
- A massive quiescent galaxy at redshift 4.658. Nat 619 (7971), pp. 716–719 (en). External Links: ISSN 1476-4687, Link, Document Cited by: §3.3.3.
- An Entropy Threshold for Strong Hα and Radio Emission in the Cores of Galaxy Clusters. ApJ 683, pp. L107. External Links: ISSN 0004-637X, Link, Document Cited by: §1.
- Constraining the properties of AGN host galaxies with spectral energy distribution modelling. A&A 576, pp. A10 (en). External Links: ISSN 0004-6361, 1432-0746, Link, Document Cited by: §4.2.1.
- Modeling the Panchromatic Spectral Energy Distributions of Galaxies. ARA&A 51 (1), pp. 393 (en). External Links: ISSN 0066-4146, Link, Document Cited by: §4.2.1.
- A z = 1.85 galaxy group in CEERS: Evolved, dustless, massive intra-halo light and a brightest group galaxy in the making. A&A 677, pp. A3. External Links: ISSN 0004-6361, Link, Document Cited by: §1, §4.2.1, §4.2.2, §4.2, §4.3.
- A Mature Galaxy Cluster at z=1.58 around the Radio Galaxy 7C1753+6311. ApJ 816, pp. 83. External Links: ISSN 0004-637X, Link, Document Cited by: §4.1, §4.1.
- The submillimetre view of massive clusters at z ∼ 0.8-1.6. MNRAS 486, pp. 3047–3058. External Links: ISSN 0035-8711, Link, Document Cited by: §1.
- The ROSAT Brightest Cluster Sample - III. Optical spectra of the central cluster galaxies. MNRAS 306, pp. 857–896. External Links: ISSN 0035-8711, Link, Document Cited by: §1.
- How does our choice of observable influence our estimation of the centre of a galaxy cluster? Insights from cosmological simulations. MNRAS 456, pp. 2566–2575. External Links: ISSN 0035-8711, Link, Document Cited by: §1.
- The hierarchical formation of the brightest cluster galaxies. MNRAS 375, pp. 2–14. External Links: ISSN 0035-8711, Link, Document Cited by: §1.
- Overview of the DESI Legacy Imaging Surveys. AJ 157, pp. 168. External Links: ISSN 0004-6256, Link, Document Cited by: §2.1.
- The flux ratio of the [N II] λλ 6548, 6583 å lines in sample of Active Galactic Nuclei Type 2. AdSpR 71, pp. 1219–1226. External Links: ISSN 0273-1177, Link, Document Cited by: §3.2.
- Distant Cluster Hunting. II. A Comparison of X-Ray and Optical Cluster Detection Techniques and Catalogs from the ROSAT Optical X-Ray Survey. ApJ 569, pp. 689–719. External Links: ISSN 0004-637X, Link, Document Cited by: §4.1.
- Cosmic metallicity evolution of Active Galactic Nuclei: implications for optical diagnostic diagrams. MNRAS 527 (3), pp. 8193–8212 (en). External Links: ISSN 0035-8711, 1365-2966, Link, Document Cited by: §4.2.1.
- Galaxy morphology in rich clusters: implications for the formation and evolution of galaxies.. ApJ 236, pp. 351–365. External Links: ISSN 0004-637X, Link, Document Cited by: §1.
- The Origin of the Brightest Cluster Galaxies. ApJ 502, pp. 141–149. External Links: ISSN 0004-637X, Link, Document Cited by: §1.
- An extreme case of galaxy and cluster co-evolution at z = 0.7. MNRAS 508, pp. 3663–3671. External Links: ISSN 0035-8711, Link, Document Cited by: §1.
- Clusters of Galaxies in the First Half of the Universe from the IRAC Shallow Survey. ApJ 684, pp. 905–932. External Links: ISSN 0004-637X, Link, Document Cited by: §4.1.
- The CatWISE Preliminary Catalog: Motions from WISE and NEOWISE Data. ApJS 247, pp. 69. External Links: ISSN 0067-0049, Link, Document Cited by: §2.1.
- Euclid preparation. III. Galaxy cluster detection in the wide photometric survey, performance and algorithm selection. A&A 627, pp. A23. External Links: Document, 1906.04707 Cited by: §1, §1.
- Euclid Quick Data Release (Q1) - Data release overview. A&A, accepted (Euclid Q1 SI), pp. arXiv:2503.15302. External Links: 2503.15302 Cited by: §2.2.2, §2.2, §4.3.
- Euclid Quick Data Release (Q1). First detections from the galaxy cluster workflow. A&A, in press (Euclid Q1 SI), https://doi.org/10.1051/0004-6361/202554937, pp. arXiv:2503.19196. External Links: Document, 2503.19196 Cited by: Figure 1, §1, §2.2, §3.1.1, §3, §4.1.
- Euclid Quick Data Release (Q1): From spectrograms to spectra: the SIR spectroscopic Processing Function. A&A, in press (Euclid Q1 SI), https://doi.org/10.1051/0004-6361/202554627, pp. arXiv:2503.15307. External Links: 2503.15307 Cited by: §2.2, §2.4.
- Euclid - ii. the vis instrument. A&A 697, pp. A2. External Links: Document, Link Cited by: §2.2.
- Euclid - iii. the nisp instrument. A&A 697, pp. A3. External Links: Document, Link Cited by: §2.2, §2.4, §3.3.2.
- Euclid - i. overview of the euclid mission. A&A 697, pp. A1. External Links: Document, Link Cited by: §1, §1, §2.2, §4.3.
- Euclid preparation. XVII. Cosmic Dawn Survey: Spitzer Space Telescope observations of the Euclid deep fields and calibration fields. A&A 658, pp. A126. External Links: Document, 2110.13928 Cited by: §2.3.
- Euclid Quick Data Release (Q1). From images to multiwavelength catalogues: the Euclid MERge Processing Function. A&A, in press (Euclid Q1 SI), https://doi.org/10.1051/0004-6361/202554586, pp. arXiv:2503.15305. External Links: 2503.15305 Cited by: §2.2.1, §2.2.
- Euclid Quick Data Release (Q1). Optical and near-infrared identification and classification of point-like X-ray selected sources. A&A, in press (Euclid Q1 SI), https://doi.org/10.1051/0004-6361/202554616, pp. arXiv:2503.15316. External Links: 2503.15316 Cited by: §2.2.2.
- Euclid preparation. I. The Euclid Wide Survey. A&A 662, pp. A112. External Links: Document, 2108.01201 Cited by: §1, §1, §4.3.
- Euclid preparation: XLV. Optical emission-line predictions of intermediate-z galaxy populations in GAEA for the Euclid Deep and Wide Surveys. A&A 689, pp. A276. External Links: Document, 2402.03436 Cited by: §4.2.1.
- Euclid preparation: LXI. Cosmic Dawn Survey: ’Pre-launch’ multiwavelength catalogues for Euclid Deep Field North and Euclid Deep Field Fornax. A&A 695, pp. A229. External Links: Document, 2408.05296 Cited by: §2.3.
- Note: https://doi.org/10.57780/esa-2853f3b Cited by: §2.2.
- Cooling Flows in Clusters of Galaxies. ARA&A 32, pp. 277–318. External Links: ISSN 0066-4146, Link, Document Cited by: §1.
- The Infrared Array Camera (IRAC) for the Spitzer Space Telescope. ApJS 154, pp. 10–17. External Links: ISSN 0067-0049, Link, Document Cited by: §2.3, §3.3.2.
- The Dark Energy Camera. ApJ 150, pp. 150. External Links: ISSN 0004-6256, Link, Document Cited by: §2.1, §2.2.1.
- Atomic data from the IRON Project: XXII. Radiative rates forforbidden transitions within the ground configuration of ions inthe carbon and oxygen isoelectronic sequences. A&AS 123 (1), pp. 159–171 (en). External Links: ISSN 0365-0138, 1286-4846, Link, Document Cited by: §3.2.
- The BPT Diagram in Cosmological Galaxy Formation Simulations: Understanding the Physics Driving Offsets at High Redshift. ApJ 926 (1), pp. 80 (en). External Links: ISSN 0004-637X, Link, Document Cited by: §4.2.1.
- Shaken Snow Globes: Kinematic Tracers of the Multiphase Condensation Cascade in Massive Galaxies, Groups, and Clusters. ApJ 854, pp. 167. External Links: ISSN 0004-637X, Link, Document Cited by: §4.2.3.
- Revealing the velocity structure of the filamentary nebula in NGC 1275 in its entirety. MNRAS 479, pp. L28–L33. External Links: ISSN 0035-8711, Link, Document Cited by: §4.2.3.
- A New Method For Galaxy Cluster Detection. I. The Algorithm. AJ 120, pp. 2148–2162. External Links: ISSN 0004-6256, Link, Document Cited by: §4.1.
- The Red-Sequence Cluster Survey. I. The Survey and Cluster Catalogs for Patches RCS 0926+37 and RCS 1327+29. ApJS 157, pp. 1–29. External Links: ISSN 0067-0049, Link, Document Cited by: §4.1.
- The High-redshift Clusters Occupied by Bent Radio AGN (COBRA) Survey: Follow-up Optical Imaging. ApJ 887 (1), pp. 50 (en). External Links: ISSN 0004-637X, Link, Document Cited by: §1.
- The Impact of Environment on the Stellar Mass-Halo Mass Relation. ApJ 860 (1), pp. 2 (en). External Links: ISSN 0004-637X, Link, Document Cited by: §1.
- The hierarchical growth of bright central galaxies and intracluster light as traced by the magnitude gap. MNRAS 538 (2), pp. 622–638 (en). External Links: ISSN 0035-8711, Link, Document Cited by: §1.
- YOLO-CL cluster detection in the Rubin/LSST DC2 simulations. A&A 695, pp. A246. External Links: Document Cited by: §1.
- Spitzer-selected z > 1.3 protocluster candidates in the LSST Deep Drilling Fields. MNRAS 527, pp. 10680–10696. External Links: ISSN 0035-8711, Link, Document Cited by: §1.
- Optical emission line nebulae in galaxy cluster cores 1: the morphological, kinematic and spectral properties of the sample. MNRAS 460, pp. 1758–1789. External Links: ISSN 0035-8711, Link, Document Cited by: §4.2.3.
- Dust-corrected Star Formation Rates of Galaxies. II. Combinations of Ultraviolet and Infrared Tracers. ApJ 741, pp. 124. External Links: ISSN 0004-637X, Link, Document Cited by: §3.2.
- Multiwavelength investigations of co-evolution of bright cluster galaxies and their host clusters. MNRAS 440, pp. 588–600. External Links: ISSN 0035-8711, Link, Document Cited by: §1.
- Beware the recent past: a bias in spectral energy distribution modelling due to bursty star formation. MNRAS 530, pp. L7–L12. External Links: ISSN 0035-8711, Link, Document Cited by: §4.2.1.
- The structure and evolution of a forming galaxy cluster at z = 1.62. MNRAS 459, pp. 387–401. External Links: ISSN 0035-8711, Link, Document Cited by: §4.1.
- Extreme AGN feedback in the MAssive Cluster Survey: a detailed study of X-ray cavities at z>0.3. MNRAS 421, pp. 1360–1384. External Links: ISSN 0035-8711, Link, Document Cited by: §1.
- Evidence of Runaway Gas Cooling in the Absence of Supermassive Black Hole Feedback at the Epoch of Cluster Formation. ApJ 898, pp. L50. External Links: ISSN 0004-637X, Link, Document Cited by: §1, §4.2.3.
- The Dark Energy Camera (DECam). arXiv e-prints, arXiv:0810.3600. External Links: Document Cited by: §2.2.1.
- The Infrared Spectrograph (IRS) on the Spitzer Space Telescope. ApJS 154, pp. 18–24. External Links: ISSN 0067-0049, Link, Document Cited by: §2.3.
- Long-slit spectroscopy of gas in the cores of X-ray luminous clusters.. ApJS 59, pp. 447–498. External Links: ISSN 0067-0049, Link, Document Cited by: §1.
- Discovering Large-scale Structure at 2 < z < 5 in the C3VO Survey. ApJ 980, pp. 155. External Links: ISSN 0004-637X, Link, Document Cited by: §1.
- LSST: From Science Drivers to Reference Design and Anticipated Data Products. ApJ 873, pp. 111. External Links: ISSN 0004-637X, Link, Document Cited by: §1.
- The dependence of star formation history and internal structure on stellar mass for 105 low-redshift galaxies. MNRAS 341, pp. 54–69. External Links: ISSN 0035-8711, Link, Document Cited by: §1.
- Effect of Local Environment and Stellar Mass on Galaxy Quenching and Morphology at 0.5 < z < 2.0. ApJ 847 (2), pp. 134 (en). External Links: Link, Document Cited by: §1, §3.1.
- On the mass function of star clusters. MNRAS 336, pp. 1188–1194. External Links: ISSN 0035-8711, Link, Document Cited by: §1, §3.3.3.
- A SINFONI view of flies in the Spiderweb: a galaxy cluster in the making. MNRAS 415, pp. 2245–2256. External Links: ISSN 0035-8711, Link, Document Cited by: §1, §4.2.1, §4.3.
- How to Measure Galaxy Star Formation Histories. II. Nonparametric Models. ApJ 876, pp. 3. External Links: ISSN 0004-637X, Link, Document Cited by: §4.2.1.
- Persistence of the colour-density relation and efficient environmental quenching to z ∼ 1.4. MNRAS 490, pp. 1231–1254. External Links: ISSN 0035-8711, Link, Document Cited by: §1.
- Information criteria for astrophysical model selection. MNRAS 377 (1), pp. L74–L78 (en). External Links: ISSN 1745-3933, 1745-3925 Cited by: §3.2.
- The importance of major mergers in the build up of stellar mass in brightest cluster galaxies at z = 1. MNRAS 433, pp. 825–837. External Links: ISSN 0035-8711, Link, Document Cited by: §1.
- Evidence for significant growth in the stellar mass of brightest cluster galaxies over the past 10 billion years. MNRAS 427, pp. 550–568. External Links: ISSN 0035-8711, Link, Document Cited by: §1.
- SWIRE: The SIRTF Wide-Area Infrared Extragalactic Survey. PASP 115, pp. 897–927. External Links: ISSN 0004-6280, Link, Document Cited by: §2.3.
- First Insights into the Spitzer Wide-Area Infrared Extragalactic Legacy Survey (SWIRE) Galaxy Populations. ApJS 154, pp. 54–59. External Links: ISSN 0067-0049, Link, Document Cited by: §2.3.
- Optical substructure and BCG offsets of Sunyaev-Zel’dovich and X-ray-selected galaxy clusters. MNRAS 478, pp. 5473–5490. External Links: ISSN 0035-8711, Link, Document Cited by: §1.
- THE FORMATION OF MASSIVE CLUSTER GALAXIES. ApJ 720 (1), pp. 284–298 (en). External Links: ISSN 0004-637X, 1538-4357, Link, Document Cited by: §3.1.
- EzGal: A Flexible Interface for Stellar Population Synthesis Models. PASP 124, pp. 606. External Links: ISSN 0004-6280, Link, Document Cited by: §4.2.1.
- The CatWISE2020 Catalog. ApJS 253 (1), pp. 8. External Links: Document, 2012.13084 Cited by: §2.1.
- Spitzer Planck Herschel Infrared Cluster (SPHerIC) survey: Candidate galaxy clusters at 1.3 < z < 3 selected by high star-formation rate. A&A 620, pp. A198. External Links: ISSN 0004-6361, Link, Document Cited by: §1.
- AMICO galaxy clusters in KiDS-DR3: sample properties and selection function. MNRAS 485, pp. 498–512. External Links: ISSN 0035-8711, Link, Document Cited by: §1.
- A Massive, Cooling-Flow-Induced Starburst in the Core of a Highly Luminous Galaxy Cluster. Nat 488 (7411), pp. 349–352. Note: Comment: 11 pages, 3 figures, 1 table. Supplemental material contains 15 additional pages. Published in Nature External Links: ISSN 0028-0836, 1476-4687, Link, Document Cited by: §4.2.3.
- Star-forming Brightest Cluster Galaxies at 0.25 < z < 1.25: A Transitioning Fuel Supply. ApJ 817, pp. 86. External Links: ISSN 0004-637X, Link, Document Cited by: §1, §1.
- An HST/WFC3-UVIS View of the Starburst in the Cool Core of the Phoenix Cluster. ApJ 765, pp. L37. External Links: ISSN 0004-637X, Link, Document Cited by: §4.2.3.
- On the Origin of the Extended Hα Filaments in Cooling Flow Clusters. ApJ 721, pp. 1262–1283. External Links: ISSN 0004-637X, Link, Document Cited by: §4.2.3.
- The heating of gas in a galaxy cluster by X-ray cavities and large-scale shock fronts. Nat 433 (7021), pp. 45–47 (en). External Links: ISSN 1476-4687, Link, Document Cited by: §1.
- Heating Hot Atmospheres with Active Galactic Nuclei. ARA&A 45 (1), pp. 117–175. Note: Comment: 60 pages, 12 figures, to appear in 1997 Annual Reviews of A&A. This version supersedes the April 2007 version in Reviews in Advance (references and minor corrections were added), and is similar to the one scheduled to appear in Volume 45 of ARAA External Links: ISSN 0066-4146, 1545-4282, Link, Document Cited by: §1.
- Mechanical Feedback from Active Galactic Nuclei in Galaxies, Groups, and Clusters. NJPh 14 (5), pp. 055023. Note: Comment: Accepted for publication in New Journal of Physics Focus Issue on Clusters of Galaxies External Links: ISSN 1367-2630, Link, Document Cited by: §1.
- Star formation in cooling flows in clusters of galaxies. AJ 98, pp. 2018–2043. External Links: ISSN 0004-6256, Link, Document Cited by: §1.
- The Spiderweb Galaxy: A Forming Massive Cluster Galaxy at z ~ 2. ApJ 650, pp. L29–L32. External Links: ISSN 0004-637X, Link, Document Cited by: §1, §4.2.2, §4.2.
- The C4 Clustering Algorithm: Clusters of Galaxies in the Sloan Digital Sky Survey. AJ 130 (3), pp. 968–1001 (en). External Links: ISSN 0004-6256, Link, Document Cited by: §4.1.
- A massive core for a cluster of galaxies at a redshift of 4.3. Nat 556, pp. 469–472. External Links: ISSN 0028-0836, Link, Document Cited by: §1, §4.2.1, §4.2.2, §4.2.2, §4.2, §4.3.
- Velocity structure functions in multiphase turbulence: interpreting kinematics of Hα filaments in cool-core clusters. MNRAS 510, pp. 2327–2343. External Links: ISSN 0035-8711, Link, Document Cited by: §4.2.3.
- The growth of brightest cluster galaxies in the TNG300 simulation: dissecting the contributions from mergers and in situ star formation. MNRAS 521, pp. 800–817. External Links: ISSN 0035-8711, Link, Document Cited by: §1.
- Reconstructing Star Formation Histories of High-redshift Galaxies: A Comparison of Resolved Parametric and Nonparametric Models. ApJ 983, pp. 181. External Links: ISSN 0004-637X, Link, Document Cited by: §4.2.1.
- Stellar mass function of cluster galaxies at z ~ 1.5: evidence for reduced quenching efficiency at high redshift. A&A 592, pp. A161. External Links: ISSN 0004-6361, Link, Document Cited by: §4.1.
- The H star formation main sequence in cluster and field galaxies at z 1.6. MNRAS 499 (3), pp. 3061–3070. External Links: Document, 2009.07861 Cited by: §1.
- Spectroscopic Confirmation of the Rich z = 1.80 Galaxy Cluster JKCS 041 using the WFC3 Grism: Environmental Trends in the Ages and Structure of Quiescent Galaxies. ApJ 788, pp. 51. External Links: ISSN 0004-637X, Link, Document Cited by: §1, §4.1.
- HST Grism Confirmation of 16 Structures at 1.4 < z < 2.8 from the Clusters Around Radio-Loud AGN (CARLA) Survey. ApJ 859, pp. 38. External Links: ISSN 0004-637X, Link, Document Cited by: §4.1.
- BASS. XXIV. The BASS DR2 Spectroscopic Line Measurements and AGN Demographics. ApJS 261 (1), pp. 4 (en). External Links: ISSN 0067-0049, 1538-4365, Link, Document Cited by: §4.2.1.
- An Observational Link between AGN Eddington Ratio and [N ii]λ6583/Hα at 0.6 < z < 1.7. ApJ 880 (2), pp. 112 (en). External Links: ISSN 0004-637X, 1538-4357, Link, Document Cited by: §4.2.1.
- Galaxy And Mass Assembly (GAMA): testing galaxy formation models through the most massive galaxies in the Universe. MNRAS 440, pp. 762–775. External Links: ISSN 0035-8711, Link, Document Cited by: §1.
- Ubiquitous cold and massive filaments in cool core clusters. A&A 631, pp. A22. External Links: ISSN 0004-6361, Link, Document Cited by: §4.2.3.
- Gas condensation in brightest group galaxies unveiled with MUSE. Morphology and kinematics of the ionized gas. A&A 666, pp. A94. External Links: ISSN 0004-6361, Link, Document Cited by: §4.2.3.
- The evolution of brightest cluster galaxies in the nearby Universe II: The star-formation activity and the stellar mass from spectral energy distribution. MNRAS 512, pp. 2758–2776. External Links: ISSN 0035-8711, Link, Document Cited by: §1.
- Systematic Identification of LAEs for Visible Exploration and Reionization Research Using Subaru HSC (SILVERRUSH). I. Program strategy and clustering properties of ∼2000 Lyα emitters at z = 6-7 over the 0.3-0.5 Gpc2 survey area. PASJ 70, pp. S13. External Links: ISSN 0004-6264, Link, Document Cited by: §1.
- A Spitzer-selected Galaxy Cluster at z = 1.62. ApJ 716, pp. 1503–1513. External Links: ISSN 0004-637X, Link, Document Cited by: §3.1.
- Detailed Structural Decomposition of Galaxy Images. AJ 124, pp. 266–293. External Links: ISSN 0004-6256, Link, Document Cited by: §3.3.2.
- Detailed Decomposition of Galaxy Images. II. Beyond Axisymmetric Models. AJ 139 (6), pp. 2097–2129 (en). External Links: ISSN 0004-6256, Link, Document Cited by: §3.3.2.
- Mass and Environment as Drivers of Galaxy Evolution in SDSS and zCOSMOS and the Origin of the Schechter Function. ApJ 721, pp. 193–221. External Links: ISSN 0004-637X, Link, Document Cited by: §1, §1, §3.1.
- Mass and Environment as Drivers of Galaxy Evolution. II. The Quenching of Satellite Galaxies as the Origin of Environmental Effects. ApJ 757, pp. 4. External Links: ISSN 0004-637X, Link, Document Cited by: §1, §1.
- The XXL Survey. I. Scientific motivations - XMM-Newton observing plan - Follow-up observations and simulation programme. A&A 592, pp. A1. External Links: ISSN 0004-6361, Link, Document Cited by: §1, §1.
- The Evolution of the Quenching of Star Formation in Cluster Galaxies since z ∼ 1. ApJ 876, pp. 40. External Links: ISSN 0004-637X, Link, Document Cited by: §1.
- Planck 2018 results. VI. Cosmological parameters. A&A 641, pp. A6 (en). External Links: ISSN 0004-6361, Link, Document Cited by: §1, §3.3.3.
- The Star Formation Histories of Galaxies in Distant Clusters. ApJ 518, pp. 576–593. External Links: ISSN 0004-637X, Link, Document Cited by: §1.
- PypeIt: The Python Spectroscopic Data Reduction Pipeline. JOSS 5, pp. 2308. External Links: Link, Document Cited by: §2.4.
- Tracing the Star-formation-Density Relation to z ~ 2. ApJ 744, pp. 88. External Links: ISSN 0004-637X, Link, Document Cited by: §4.1.
- The Feedback-regulated Growth of Black Holes and Bulges through Gas Accretion and Starbursts in Cluster Central Dominant Galaxies. ApJ 652, pp. 216–231. External Links: ISSN 0004-637X, Link, Document Cited by: §1.
- The Regulation of Cooling and Star Formation in Luminous Galaxies by Active Galactic Nucleus Feedback and the Cooling-Time/Entropy Threshold for the Onset of Star Formation. ApJ 687, pp. 899–918. External Links: ISSN 0004-637X, Link, Document Cited by: §1.
- Galaxies at the edges: A complete census of the MACS J0416.1─2403 cluster. A&A 701, pp. A38. External Links: ISSN 0004-6361, Link, Document Cited by: §1.
- Galaxy mass, cluster-centric distance and secular evolution: their role in the evolution of galaxies in clusters in the last 10 Gyr. A&A 543, pp. A19. External Links: ISSN 0004-6361, Link, Document Cited by: §1.
- The Relation between Cool Cluster Cores and Herschel-detected Star Formation in Brightest Cluster Galaxies. ApJ 747, pp. 29. External Links: ISSN 0004-637X, Link, Document Cited by: §1.
- Directly imaging the cooling flow in the Phoenix cluster. Nat 638, pp. 360–364. External Links: ISSN 0028-0836, Link, Document Cited by: §4.2.3.
- The Young and the Wild: What Happens to Protoclusters Forming at Redshift z ≈ 4?. ApJ 950, pp. 191. External Links: ISSN 0004-637X, Link, Document Cited by: §1, §4.2.2.
- Rapid early coeval star formation and assembly of the most-massive galaxies in the Universe. MNRAS 493, pp. 4607–4621. External Links: Link, Document Cited by: §1, §4.2.2, §4.2.2, §4.2.2.
- The Multiband Imaging Photometer for Spitzer (MIPS). ApJS 154, pp. 25–29. External Links: ISSN 0067-0049, Link, Document Cited by: §2.3, §2.3.
- Galaxy formation and evolution science in the era of the Large Synoptic Survey Telescope. NatRP 1 (7), pp. 450–462 (en). External Links: ISSN 2522-5820, Link, Document Cited by: §1.
- The SVO Filter Profile Service. In XIV.0 Scientific Meeting (virtual) of the Spanish Astronomical Society, pp. 182. Cited by: §2.2.1.
- Photometric segregation of dwarf and giant FGK stars using the SVO Filter Profile Service and photometric tools. A&A 689, pp. A93 (en). External Links: ISSN 0004-6361, Link, Document Cited by: §2.2.1.
- SVO Filter Profile Service Version 1.0. Note: IVOA Working Draft 15 October 2012https://doi.org/10.5479/ADS/bib/2012ivoa.rept.1015R External Links: Document Cited by: §2.2.1.
- Optical and near-infrared observations of the SPT2349-56 proto-cluster core at z = 4.3. MNRAS 502, pp. 1797–1815. External Links: ISSN 0035-8711, Link, Document Cited by: §4.3.
- Optical Spectroscopy and Velocity Dispersions of Galaxy Clusters from the SPT-SZ Survey. ApJ 792, pp. 45. External Links: ISSN 0004-637X, Link, Document Cited by: §4.2.1.
- redMaPPer. I. Algorithm and SDSS DR8 Catalog. ApJ 785, pp. 104. External Links: ISSN 0004-637X, Link, Document Cited by: §4.1.
- GALEX-SDSS-WISE Legacy Catalog (GSWLC): Star Formation Rates, Stellar Masses, and Dust Attenuations of 700,000 Low-redshift Galaxies. ApJS 227, pp. 2. External Links: ISSN 0067-0049, Link, Document Cited by: §4.2.1.
- Next generation cosmology: constraints from the Euclid galaxy cluster survey. MNRAS 459, pp. 1764–1780. External Links: ISSN 0035-8711, Link, Document Cited by: §1.
- Rapid Coeval Black Hole and Host Galaxy Growth in MRC 1138-262: The Hungry Spider. ApJ 755, pp. 146. External Links: ISSN 0004-637X, Link, Document Cited by: §4.2.1.
- Wide-Field InfrarRed Survey Telescope-Astrophysics Focused Telescope Assets WFIRST-AFTA 2015 Report. arXiv e-prints, arXiv:1503.03757. External Links: Document Cited by: §1.
- Simulations of the formation, evolution and clustering of galaxies and quasars. Nat 435, pp. 629–636. External Links: ISSN 0028-0836, Link, Document Cited by: §1.
- IDCS J1426.5+3508: Discovery of a Massive, Infrared-selected Galaxy Cluster at z = 1.75. ApJ 753, pp. 164. External Links: ISSN 0004-637X, Link, Document Cited by: §3.2.
- Theoretical values for the [OIII] 5007/4959 line-intensity ratio and homologous cases. MNRAS 312, pp. 813–816. External Links: ISSN 0035-8711, Link, Document Cited by: §3.2.
- The Red Sequence at Birth in the Galaxy Cluster Cl J1449+0856 at z = 2. ApJ 833, pp. L20 (en). External Links: Link, Document Cited by: §3.1.
- Galaxy populations in the most distant SPT-SZ clusters. I. Environmental quenching in massive clusters at 1.4 ≲ z ≲ 1.7. A&A 622, pp. A117 (en). External Links: Link, Document Cited by: §1, §4.1.
- Bright [C II]158 μm Streamers as a Beacon for Giant Galaxy Formation in SPT2349-56 at z = 4.3. ApJ 998, pp. 191. External Links: ISSN 0004-637X, Link, Document Cited by: §4.2.2.
- The Massive and Distant Clusters of WISE Survey 2: Second Data Release. ApJ 976, pp. 186. External Links: ISSN 0004-637X, Link, Document Cited by: §1, §2.1, §4.1.
- The Massive and Distant Clusters of WISE Survey 2: Equatorial First Data Release. ApJ 967, pp. 123. External Links: ISSN 0004-637X, Link, Document Cited by: §1, §1, §2.1.
- COSMOS-Web galaxy groups: Evolution of red sequence and quiescent galaxy fraction. A&A 707, pp. A87. External Links: ISSN 0004-6361, Link, Document Cited by: §1.
- GOLDRUSH. III. A systematic search for protoclusters at z ∼ 4 based on the >100 deg2 area. PASJ 70, pp. S12. External Links: ISSN 0004-6264, Link, Document Cited by: §1.
- Reversal of Fortune: Confirmation of an Increasing Star Formation-Density Relation in a Cluster at z = 1.62. ApJ 719, pp. L126–L129. External Links: ISSN 0004-637X, Link, Document Cited by: §1.
- The Massive and Distant Clusters of WISE Survey 2: A Stacking Analysis Investigating the Evolution of Star Formation Rates and Stellar Masses in Groups and Clusters. ApJ 972, pp. 27. External Links: ISSN 0004-637X, Link, Document Cited by: §1.
- Multiwavelength radio observations of a brightest cluster galaxy at z = 1.71: detection of a modest active galactic nucleus and evidence for extended star formation. MNRAS 487 (1), pp. 1210 (en). External Links: Link, Document Cited by: §4.2.3.
- The XMM-LSS survey. First high redshift galaxy clusters: Relaxed and collapsing systems. A&A 423, pp. 75–85. External Links: ISSN 0004-6361, Link, Document Cited by: §4.1.
- The GOGREEN Survey: A deep stellar mass function of cluster galaxies at 1.0 < z < 1.4 and the complex nature of satellite quenching. A&A 638, pp. A112 (en). External Links: ISSN 0004-6361, Link, Document Cited by: §1, §3.1.1.
- Composite Quasar Spectra from the Sloan Digital Sky Survey. ApJ 122, pp. 549–564. External Links: ISSN 0004-6256, Link, Document Cited by: §3.3.3.
- Chandra Cluster Cosmology Project III: Cosmological Parameter Constraints. ApJ 692 (2), pp. 1060–1074. External Links: Document, 0812.2720 Cited by: §1.
- Introducing the Illustris Project: simulating the coevolution of dark and visible matter in the Universe. MNRAS 444, pp. 1518–1547. External Links: ISSN 0035-8711, Link, Document Cited by: §1.
- Star formation and AGN activity in SDSS cluster galaxies. MNRAS 404, pp. 1231–1246. External Links: ISSN 0035-8711, Link, Document Cited by: §1.
- The Star Formation History of BCGs to z = 1.8 from the SpARCS/SWIRE Survey: Evidence for Significant In Situ Star Formation at High Redshift. ApJ 814, pp. 96. External Links: ISSN 0004-637X, Link, Document Cited by: §1.
- An Extreme Starburst in the Core of a Rich Galaxy Cluster at z = 1.7. ApJ 809, pp. 173. External Links: ISSN 0004-637X, Link, Document Cited by: §1, §3.1, §3.2, §4.2.1, §4.2.2, §4.2.3, §4.2.
- The Spitzer Space Telescope Mission. ApJS 154, pp. 1–9. External Links: ISSN 0067-0049, Link, Document Cited by: §2.3.
- Spectroscopic confirmation of a mature galaxy cluster at a redshift of 2. Nat 577, pp. 39–41. External Links: ISSN 0028-0836, Link, Document Cited by: §3.1.
- Clusters of Galaxies at 1 < z < 2 : The Spitzer Adaptation of the Red-Sequence Cluster Survey. arXiv e-prints, astro-ph:0604289, pp. . External Links: Document Cited by: §1, §4.1.
- Spectroscopic Confirmation of a Massive Red-Sequence-Selected Galaxy Cluster at z = 1.34 in the SpARCS-South Cluster Survey. ApJ 698, pp. 1943–1950. External Links: ISSN 0004-637X, Link, Document Cited by: §1, §4.1.
- Dependence of galaxy quenching on halo mass and distance from its centre. MNRAS 428, pp. 3306–3326. External Links: ISSN 0035-8711, Link, Document Cited by: §4.1.
- The Wide-field Infrared Survey Explorer (WISE): Mission Description and Initial On-orbit Performance. AJ 140, pp. 1868–1881. External Links: ISSN 0004-6256, Link, Document Cited by: §1.
- Galaxy Clusters around Radio-loud Active Galactic Nuclei at 1.3 < z < 3.2 as Seen by Spitzer. ApJ 769, pp. 79. External Links: ISSN 0004-637X, Link, Document Cited by: §1.
- MAMMOTH-MOSFIRE: Environmental Effects on Galaxy Interstellar Medium at z ∼ 2. ApJ 993 (2), pp. 231 (en). External Links: ISSN 0004-637X, 1538-4357, Link, Document Cited by: §4.2.1.
- Turbulent Heating in Galaxy Clusters Brightest in X-rays. Nat 515 (7525), pp. 85–87. Note: Comment: published in Nature External Links: ISSN 0028-0836, 1476-4687, Link, Document Cited by: §1.
- Miscentring in galaxy clusters: dark matter to brightest cluster galaxy offsets in 10 000 Sloan Digital Sky Survey clusters. MNRAS 426, pp. 2944–2956. External Links: ISSN 0035-8711, Link, Document Cited by: §1.